# Composite likelihood estimation of stationary Gaussian processes

### v1

@@ Abstract
We develop a framework for composite likelihood inference of parametric continuous-time stationary Gaussian processes. We derive the asymptotic theory of the associated maximum composite likelihood estimator. We implement our approach on a pair of models that has been proposed to describe the random log-spot variance of financial asset returns. A simulation study shows that it delivers good performance in these settings and improves upon a method-of-moments estimation. In an application, we inspect the dynamic of an intraday measure of spot variance computed with high-frequency data from the cryptocurrency market. The empirical evidence supports a mechanism, where the short- and long-term correlation structure of stochastic volatility are decoupled in order to capture its properties at different time scales.
JEL Classification: C10; C80.
@@ 1 Introduction
The search for the most accurate description of the dynamic of time-varying volatility continues to permeate financial economics. On the one hand, it has long been recognized that realized variance is highly persistent and exhibits an autocorrelation function (acf) that perishes so slowly that it can best be approximated by a stochastic process featuring long memory (see, e.g., Andersen, Bollerslev, Diebold, and Labys, 2003; Comte and Renault, 1998; Corsi, 2009).
@@
On the other hand, in recent years it has been suggested that the sample path of stochastic volatility may be more vibrant than implied by a standard Brownian motion. This idea was put forward by Gatheral, Jaisson, and Rosenbaum (2018), who observed that the curvature of the implied volatility surface for short-term at-the-money options is consistent with stochastic volatility being governed by a fractional Brownian motion (fBm) with a Hurst exponent less than a half. Since then, a sequence of follow-up papers studied various statistical procedures for gauging the roughness index of volatility, both under the physical and risk-neutral probability measure, largely confirming that “volatility is rough” (e.g., Bennedsen, 2020; Bolko, Christensen, Pakkanen, and Veliyev, 2023; Chong and Todorov, 2023; Wang, Xiao, and Yu, 2023). [^2] A comprehensive overview of the literature is available at the Rough Volatility Network’s website: https://sites.google.com/site/roughvol/home/rough-volatility-literature.
A workhorse in this literature is the fractional Ornstein-Uhlenbeck (fOU) process, where the standard Brownian motion in the Gaussian Ornstein-Uhlenbeck process is replaced with an fBM as impetus. This model is parametric, so in principle we can do full maximum likelihood estimation (MLE) of it (Wang, Xiao, Yu, and Zhang, 2023).
In practice, MLE of the fOU process is challenging for several reasons. First, the model is generally non-Markovian. Second, in the volatility literature the state variable in the fOU represents the point-in-time log-variance of an asset price process. Previous work has employed the daily log-realized variance as an observable surrogate. However, roughness is a sample path property that concerns the behavior of a stochastic process at very short time scales. This suggests that estimation of the fOU process should be based on much more localized measure of log-variance. In this paper, we recover the log-spot variance at the intraday horizon.
@@
However, with a long sample of a discretely observed log-spot variance process even Gaussian MLE can be prohibitive—except in special cases—because the calculation of the determinant and inversion of the covariance matrix are computationally intensive, when the number of observations is moderate-to-large, e.g. in excess of 1,000. To circumvent this issue, we propose a composite likelihood estimator for parametric stationary Gaussian processes, which nests the fOU. The composite likelihood approach—introduced in Lindsay (1988)—reduces the complexity of full MLE by only including lower-dimensional sub-models in the objective function. This effectively allows it to operate on arbitrarily large samples with minimal extra computational cost. Furthermore, while it loses asymptotic efficiency compared to MLE, in practice composite likelihood can be preferred when the full likelihood is impractical or difficult to evaluate. Moreover, in finite samples the loss of efficiency may not be too severe, since the composite likelihood function can be smoother than the full likelihood surface and, hence, much more convenient to navigate and optimize. The theoretical properties of the composite likelihood estimator have been examined in various settings, see, e.g., Cox and Reid (2004), Davis and Yau (2011) and Varin, Reid, and Firth (2011). It has also been applied to many fields, including finance (Bennedsen, Lunde, Shephard, and Veraart, 2023; Pakel, Shephard, Sheppard, and Engle, 2020).
In previous work, the fOU process has been examined with a method-of-moments estimator (MME), e.g. Bolko, Christensen, Pakkanen, and Veliyev (2023) and Wang, Xiao, and Yu (2023). [^3] Notable exceptions are Fukasawa, Takabatake, and Westphal (2022) and Shi, Yu, and Zhang (2023), who develop Whittle-type approximate likelihood estimation of the fOU model. The advantages of the maximum composite likelihood estimator (MCLE) over the MME are at least threefold. First, the MME for the Hurst parameter often relies on in-fill asymptotic theory to derive consistency, while our composite likelihood theory is derived within a long-span setting. The latter is appropriate, when the process is sampled discretely on a fixed equidistant partition over an expanding horizon. Second, our simulation results document that MCLE is more accurate than the MME. Thirdly, MCLE extracts the entire parameter vector in a single step rather than via a two-stage approach.
A crucial weakness of the fOU process as a model for the random log-volatility of financial asset returns is that controls both the short- and long-run persistence in a single parameter, namely the Hurst exponent. However, roughness is a sample path property of a stochastic process (leading to a rapid decline in the acf at short time scales), whereas long memory is a property of the distribution function (leading to a slow decline in the acf at long time scales). As these effects are interwined in the fOU process, it is only capable of featuring either roughness or long memory. This shortcoming was in fact highlighted by Mandelbrot (1982) in the context of the fBm. Set against this backdrop, we explore a stationary Gaussian process from a so-called Cauchy class (Gneiting and Schlather, 2004). As for the fOU process, the latter has three parameters to fit the dynamic dependence, but it reserves a separate parameter to describe the short- and long-term decay. Hence, it allows to decouple the behavior of the process at different time scales and is therefore able to account for both roughness and long memory.
@@
The most closely related paper within the composite likelihood literature is Davis and Yau (2011). They study the theoretical properties of a pairwise likelihood estimator in a standard time series setting with discrete-time linear processes—not necessarily Gaussian—where the dependence structure is determined by the decay of the coefficients in the linear filter. By contrast, our processes evolve in continuous-time. Moreover, we develop our theory in the general framework of $q$-wise composite likelihood. Even in the pairwise setting, however, our asymptotic theory extends Davis and Yau (2011) by allowing for the presence of a slowly varying function at infinity in the autocorrelation structure of the process. This is relevant outside the fOU model, such as the Brownian semistationary process with a power kernel (e.g., Barndorff-Nielsen and Schmiegel, 2009).
@@
In our empirical application, we inspect a vast high-frequency dataset from the cryptocurrency market. We compute a measure of spot variance based on an intraday realized variance estimator. We implement the MCLE procedure on this series in order to estimate both the fOU process and the Cauchy class. The results are striking in that the fOU, in agreement with recent work, suggests that the log-spot variance is rough, i.e. it is described by a Hurst parameter less than a half. Conversely, the Cauchy class strongly points toward a dynamic, where both roughness and long memory are required to describe the time-varying volatility. This confirm the findings from equity high-frequency data in related work of Bennedsen, Lunde, and Pakkanen (2022).
@@
The rest of the paper is structured as follows. In Section 2, we introduce the class of stationary Gaussian processes and composite likelihood estimation. We also derive the asymptotic behaviour of the MCLE. In Section 3, we introduce the fOU process and the Cauchy class that are the concrete parametric models we investigate more in-depth. Section 4 presents a simulation study of our estimator within this framework and documents its efficacy in small samples, also compared to a MME. Section 5 contains empirical work, where we implement the technique on a high-frequency time series of log-spot variance estimates from a number of cryptocurrency spot exchange rates. In Section 6, we conclude and point to directions for further work. An appendix contains mathematical derivations.
@@ 2.1 Stationary Gaussian processes
In this paper, we look at processes of the form
@@
where $\mu\in\mathbb{R}$, $\nu>0$, and $X=(X_{t})_{t\in\mathbb{R}}$ is a stationary Gaussian process with mean zero and unit variance, i.e. the marginal distribution of $X_{t}$ is standard normal. We denote the autocovariance function (acf) of $Y$ at lag $h$ as $\gamma_{h}=\text{cov}(Y_{t},Y_{t+h})=\nu^{2}E(X_{t}X_{t+h})=\nu^{2}\rho_{h}$, where $\rho_{h}$ is the acf of $X$ [^4] Throughout the paper, we loosely employ acf to represent both the autocovariance function and autocorrelation function. These are identical for $X$ and related through a scaling by $\nu^{2}$ for $Y$.
@@
A stationary Gaussian process is ergodic if and only Assumption 2.1 is fulfilled (Maruyama, 1949). Since we need a law of large numbers to hold for the log-composite likelihood function, the above can therefore be viewed as a minimal regularity condition.
@@ 2.2 Composite likelihood
We suppose $\Delta>0$ is a fixed time gap and let $y=(y_{1},y_{2},\dots,y_{n})^{\top}$ be a discrete realization of $n$ equidistant observations of the random vector $(Y_{\Delta},Y_{2\Delta},\dots,Y_{n\Delta})^{\top}$, where $Y$ is the stationary continuous-time process in (2.1). We assume $\theta\in\Theta$ is a finite-dimensional parameter vector that determines the distribution of $Y$, where $\Theta\subset\mathbb{R}^{p}$ is a compact set. The data-generating value of $\theta$ is denoted $\theta_{0}$. Moreover, we initially enforce that $\mu=0$ and this is known to the econometrician. In Remark 1, we elaborate further on the effect of estimating the mean.
@@
$l(\theta;y)\propto-\log|\Sigma_{n}|-y^{\top}\Sigma_{n}^{-1}y,$
and $\Sigma_{n}$ is the $n\times n$ covariance matrix of $(Y_{\Delta},Y_{2\Delta},\dots,Y_{n\Delta})^{\top}$.
The maximum likelihood estimator, $\hat{\theta}_{\text{MLE}}$, is consistent, asymptotically efficient, and follows a limiting normal distribution under standard regularity conditions (e.g., Wooldridge, 1994).
However, even in the Gaussian setting, moderate values of $n$ can render numerical optimization of the log-likelihood function infeasible, because the complexity of computing the determinant and inverting the covariance matrix grows very rapidly. For example, deploying an algorithm such as the Cholesky factorization or LU decomposition results in a computational budget of $O(n^{3})$. Less naive approaches exploit the fact that the covariance matrix of a stationary process observed on an equidistant partition has a Toeplitz structure (e.g., Levinson, 1946; Durbin, 1960; Trench, 1964), yielding a faster calculation speed of $O(n^{2})$. An extra refinement has been achieved with so-called “superfast” algorithms that operate on Toeplitz matrices using the fast Fourier transform (see Brent, Gustavson, and Yun, 1980). The latter exhibit near-linear growth $O(n\log^{2}(n))$ but suffer from numerical instability in practice (see Stewart, 2003). Moreover, the scaling of these algorithms are still inferior to the linear rate $O(n)$. This poses serious issues in a high-frequency setting, where $n$ is often prohibitively large.
Set against this backdrop, we exploit the composite likelihood framework of Lindsay (1988), building on the earlier concept of pseudolikelihood from Besag (1974) in the spatial setting, see also the survey by Varin, Reid, and Firth (2011). To describe the idea, assume for the moment that $y$ is a sample of length $n$ of a continuous random variable $Y$ defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. [^5] In our setting, the observations are a discrete realization of a continuous-time process. In general, the composite likelihood estimator maximizes a weighted product of likelihoods of marginal or conditional events. We let $f$ denote the density of $Y$ and suppose that $(\mathcal{A}_{1},\dots,\mathcal{A}_{M})$ is a collection of events, $A_{m}\in\mathcal{F}$, with likelihood $L_{m}(\theta;y)\propto f(y\in\mathcal{A}_{m};\theta)$. The composite likelihood is then defined as
@@
The composite log-likelihood function is, in general, not proportional to the full likelihood (and therefore not always a genuine likelihood), so the model is misspecified (see, e.g., White, 1982). However, this is not an example, where the data-generating probability measure is completely unrelated to the assumed parametric distribution. Here, the marginal (or conditional) densities included in (2.4) are extracted from the true model and correctly specified. Hence, the score of $cl(\theta;y)$ satisfies the first Bartlett identity and forms a collection of unbiased estimating equations. As we show below, it follows by a law of large numbers that $\hat{\theta}_{\mathrm{MCLE}}$ is consistent for the true parameter value, $\theta_{0}$, rather than some pseudo-parameter minimizing the Kullback-Leibler divergence between the true and assumed model. Moreover, if the memory in the process is not too strong, $\hat{\theta}_{\mathrm{MCLE}}$ is also asymptotically normal.
2.3 Asymptotic theory
@@
In the present paper, we establish an asymptotic theory for a $q$-wise composite likelihood estimator, where the composite likelihood function is formed as an equal weighted product of marginal events defined via the selection of tuples of up to length $q$. It nests both the independence likelihood and the pairwise likelihood as special cases. We take $K\in\mathbb{N}$ fixed and let $Q$ be a collection of $K$ $q^{j}$-tuples of natural numbers, where $q^{j}\in\mathbb{N}$, i.e. tuples of the form $k^{j}=(0,k_{2}^{j},\dots,k_{q^{j}}^{j})\in\{0\}\times\mathbb{N}^{q^{j}-1}$ for $j=1,\dots,K$, indexing which observations to add. Here, we use the convention that $\mathbb{N}^{0}=\emptyset$. To ease notation, we set $q=\max_{j=1,\dots,K}q^{j}$. Moreover, we suppose without loss of generality that the indices are increasing and that $\max_{j=1,\dots,K}k_{q^{j}}^{j}<n$. For $k^{j}\in Q$, we let $f_{k^{j}}(y_{i}^{k^{j}};\theta)=f_{k^{j}}(y_{i},y_{i+k_{2}^{j}},\dots,y_{i+k_{q^{j}}^{j}};\theta)$ denote the density of $Y_{i}^{k^{j}}=(Y_{i},Y_{i+k_{2}^{j}},\dots,Y_{i+k_{q^{j}}^{j}})$, which by stationary is independent of $i$.
@@
To prove our asymptotic theory, we need a identification assumption to ensure that the log-composite likelihood is uniquely maximized at $\theta_{0}$:

### v2

# To be or not to be:

@@ Abstract
We develop a framework for composite likelihood estimation of parametric continuous-time stationary Gaussian processes. We derive the asymptotic theory of the associated maximum composite likelihood estimator. We implement our approach on a pair of models that have been proposed to describe the random log-spot variance of financial asset returns. A simulation study shows that it delivers good performance in these settings and improves upon a method-of-moments estimation. In an empirical investigation, we inspect the dynamic of an intraday measure of the spot log-realized variance computed with high-frequency data from the cryptocurrency market. The evidence supports a mechanism, where the short- and long-term correlation structure of stochastic volatility are decoupled in order to capture its properties at different time scales. This is further backed by an analysis of the associated spot log-trading volume.
JEL Classification: C10; C58; C80.
@@ 1 Introduction
The search for the best description of the dynamic of time-varying volatility continues to permeate financial economics. On the one hand, it has long been recognized that realized variance (RV) is highly persistent and exhibits an autocorrelation function (ACF) that perishes so slowly that it can be approximated by a stochastic process featuring long memory (see, e.g., Comte and Renault, 1998; Andersen, Bollerslev, Diebold, and Labys, 2003; Corsi, 2009). On the other hand, in recent years it has been suggested that the sample path of volatility is more vibrant than those implied by a standard Brownian motion (sBm), and that it may instead be generated by a fractional Brownian motion (fBm) with a Hurst exponent less than a half; the index of the sBm. Gatheral, Jaisson, and Rosenbaum (2018) observe that the shape of the implied volatility surface for short-term at-the-money options is consistent with stochastic volatility being governed by such a fBm. They reinforce this finding with variogram regressions for daily log-RV series from a number of leading stock indexes, which further support their claim that “volatility is rough.” Since then, a sequence of follow-up papers have studied various statistical estimation procedures for gauging the roughness of volatility, both under the physical and risk-neutral probability measure, largely confirming this view (e.g., Fukasawa, Takabatake, and Westphal, 2022; Bolko, Christensen, Pakkanen, and Veliyev, 2023; Chong and Todorov, 2023; Wang, Xiao, and Yu, 2023; Shi, Yu, and Zhang, 2024). [^2] A more comprehensive overview of the extant literature is available at the Rough Volatility Network’s website: https://sites.google.com/site/roughvol/home/rough-volatility-literature.
The workhorse in the recent continuous-time literature on rough volatility is the fractional Ornstein-Uhlenbeck (fOU) process, where the sBm in the Gaussian Ornstein-Uhlenbeck process is replaced with an fBm as impetus. This model is parametric, so in principle we can do maximum likelihood estimation (MLE) of it. This idea is explored in a recent prominent contribution of Wang, Xiao, Yu, and Zhang (2025). In practice, MLE of the fOU process can be challenging, however. This is because the model is generally non-Markovian, which makes it difficult to estimate, because the process can depend on its entire history. [^3] Damian and Frey (2024) study a particle filter technique to estimate roughness in spot volatility. It exploits the notion that the fBm has an infinite-dimensional Markovian representation approximable by a finite sum of Ornstein-Uhlenbeck processes driven by a single sBm. However, their approach centers around the nonstationary fBm, not the stationary fOU, and furthermore it cannot capture long memory. Thus, Gaussian MLE can be prohibitive with a large sample of such a discretely observed log-spot variance processes—except in special cases—because the calculation of the determinant and inversion of the covariance matrix are computationally intensive, when the number of observations is moderate-to-large, e.g. in excess of 1,000.
@@
To circumvent this issue, in this paper we propose a composite likelihood estimator for parametric stationary Gaussian processes, which nests the fOU as a special case. The composite likelihood approach—introduced in Lindsay (1988)—reduces the complexity of MLE by only including lower-dimensional sub-models in the criterion function. This effectively allows it to operate on arbitrarily large samples with minimal extra computational cost. We derive the asymptotic theory for the maximum composite likelihood estimator (MCLE), which turns out to be somewhat intractable, because it features a discontinuity both in the rate of convergence and the shape of the limiting distribution, as a function of the memory of the process (pursuant to Hosking, 1996). Inference can also depend on whether the location parameter is estimated or not. However, while MCLE loses some of the desirable statistical properties of MLE, in practice composite likelihood can be preferred when the full likelihood is impractical or difficult to evaluate. Moreover, in finite samples the loss of efficiency may not be too severe, since the composite likelihood function can be smoother than the full likelihood surface and, hence, much more convenient to navigate and optimize (e.g., Varin, Reid, and Firth, 2011).
@@
The theoretical properties of the composite likelihood estimator have been examined in various settings, see, e.g., Cox and Reid (2004), Davis and Yau (2011), and Varin, Reid, and Firth (2011). It has also been applied to many fields, including finance (Bennedsen, Lunde, Shephard, and Veraart, 2023; Pakel, Shephard, Sheppard, and Engle, 2020). The most closely related paper to ours within the composite likelihood literature is Davis and Yau (2011). They study the theoretical properties of a pairwise MCLE in a standard time series setting with discrete-time linear processes—not necessarily Gaussian—where the dependence structure is determined by the decay of the coefficients in the linear filter. By contrast, our processes evolve in continuous-time. Moreover, we develop our theory in the general setting of $q$-wise composite likelihood. Even in the pairwise setting, however, our framework extends Davis and Yau (2011) by allowing for the presence of a slowly varying function at infinity in the autocorrelation structure of the process. This is relevant outside the fOU model, such as the Brownian semistationary process with a power kernel (e.g., Barndorff-Nielsen and Schmiegel, 2009).
In the volatility literature, the state variable in the fOU represents the point-in-time log-variance of an asset log-price, which is latent. Previous work has employed the daily log-RV as an observable surrogate, but this is an estimator of the daily log-integrated variance, i.e. an infinite mixture of log-normal random variables, which is itself not log-normal. This loses the analytic tractability of Gaussian MLE and makes the calculation of the exact transition density a nontrivial exercise. Even worse, if it is not properly accounted for, it runs the risk of mixing theoretical properties of a spot process with empirical properties of an integrated process. Moreover, since roughness is a sample path property that concerns the behavior of a stochastic process at very short time scales, this discussion suggests that estimation of the fOU process perhaps ought to be based on more localized measure of log-variance. In this paper, we recover the log-spot variance at the intraday horizon. As a result, the sample size in our empirical application is so large that MCLE is possibly the only feasible likelihood-type procedure available.
A crucial weakness of the fOU process as a model for the random log-volatility of financial asset returns is that it controls both the short- and long-run persistence in a single parameter, namely the Hurst exponent. However, roughness is a sample path property of a stochastic process (leading to a rapid decline in the ACF at short time scales), whereas long memory is a property of the distribution function (leading to a slow decline in the ACF at long time scales). As these effects are intertwined in the fOU process, it is only capable of featuring either roughness or long memory. This shortcoming was, in fact, already highlighted by Mandelbrot (1982) in the context of the fBm. Set against this backdrop, we explore a stationary Gaussian process from a so-called Cauchy class (Gneiting and Schlather, 2004). As for the fOU process, the latter has two parameters to fit the dynamic dependence, but it reserves a separate parameter to describe the short- and long-term decay. Hence, it allows to decouple the behavior of the process at different time scales and is therefore able to account for both roughness and long memory.
In a simulation study, we examine the small sample properties of the MCLE for the fOU process and the Cauchy class, which we benchmark against a method-of-moments estimator (MME), e.g. Bolko, Christensen, Pakkanen, and Veliyev (2023) and Wang, Xiao, and Yu (2023). [^4] Fukasawa, Takabatake, and Westphal (2022) and Shi, Yu, and Zhang (2024) develop Whittle-type approximate likelihood estimation of the fOU model. The advantages of the MCLE over the MME are at least twofold. First, the MME for the Hurst parameter often relies on in-fill asymptotic theory to derive consistency, while our composite likelihood theory is derived within a long-span setting. The latter is appropriate, when the process is sampled discretely on a fixed equidistant partition over an expanding horizon. Second, MCLE extracts the entire parameter vector in a single step rather than via a two-stage approach, as done in the MME. The latter appears to lead to systematic biases in the parameter estimates. As we show, this means that MCLE is often more accurate than the MME, or at the very least heading in the right direction (toward the true parameter value). However, MCLE is not uniformly superior to MME, because it depends on the exact implementation.
@@
In our empirical application, we inspect a vast high-frequency dataset from the cryptocurrency market. The coins we study are extremely liquid, so we can sample them at a much higher frequency than what is possible for more traditional assets that are traded less often. We compute an observable measure of the latent spot variance based on an intraday RV estimator. We implement the MCLE procedure on the log-series and estimate both the fOU process and the Cauchy class. The results are striking in that the fOU process, in agreement with recent work, suggests that the log-spot variance is rough. Conversely, the Cauchy class strongly points toward a dynamic, where both roughness and long memory are required to describe the time-varying volatility. This confirms the findings from equity high-frequency data in related work of Bennedsen, Lunde, and Pakkanen (2022). The main shortcoming with this application is that the RV measure is a noisy proxy of the spot variance and the inherent sampling error can distort the inferred dynamic of the variance process. Motivated by the acclaimed volume-volatility relationship, we also conduct an investigation of the trading volume that is directly observed. In the biggest estimation for log-trading volume, the MCLE is confronted with a sample of about 13 million observations, but it converges to an optimum within a few minutes. In sum, our results for log-trading volume largely confirm those from the log-RV. We conclude that both roughness and long memory are needed to describe stochastic volatility.
@@
The rest of the paper is structured as follows. In Section 2, we introduce the class of stationary Gaussian processes and composite likelihood estimation. We also derive the asymptotic behavior of the MCLE. In Section 3, we introduce the fOU process and the Cauchy class that are the concrete parametric models we investigate more in-depth. Section 4 presents a simulation study of our estimator within this framework and documents its efficacy in small samples, also compared to a MME. Section 5 contains empirical work, where we implement the technique on a high-frequency time series of log-spot variance estimates and log-trading volume from a number of cryptocurrency spot exchange rates. In Section 6, we conclude and point to directions for further work. An appendix contains mathematical derivations, presents a method-of-moments estimator, and includes supplemental simulation analysis and empirical results.
@@ 2.1 Stationary Gaussian processes
In this paper, we consider processes of the form
@@
where $\mu\in\mathbb{R}$, $\nu>0$, and $X=(X_{t})_{t\in\mathbb{R}}$ is a stationary Gaussian process with mean zero and unit variance, i.e. the marginal distribution of $X_{t}$ is standard normal. We denote the autocovariance function (ACF) of $Y$ at lag $h$ as $\gamma_{h}=\text{cov}(Y_{t},Y_{t+h})=\nu^{2}E(X_{t}X_{t+h})=\nu^{2}\rho_{h}$, where $\rho_{h}$ is the ACF of XX. [^5] Throughout the paper, we loosely employ ACF to represent both the autocovariance function and autocorrelation function. These are identical for $X$ and related through a scaling by $\nu^{2}$ for $Y$.
@@
A stationary Gaussian process is ergodic if and only if Assumption 2.1 is fulfilled (Maruyama, 1949). Since we need a law of large numbers to hold for the log-composite likelihood function, the above can therefore be viewed as a minimal regularity condition.
@@ 2.2 Composite likelihood
We suppose $\Delta>0$ is a fixed time gap and let $y=(y_{1},y_{2},\dots,y_{n})^{\top}$ be a discrete realization of $n$ equidistant observations of the random vector $Y_{n}^{\Delta}=(Y_{\Delta},Y_{2\Delta},\dots,Y_{n\Delta})^{\top}$, where $Y$ is the stationary continuous-time process in (2.1). We assume $\theta\in\Theta$ is a finite-dimensional parameter vector that determines the distribution of $Y$, where $\Theta\subset\mathbb{R}^{p}$ is a compact set. The data-generating value of $\theta$ is denoted $\theta_{0}$. Moreover, we initially enforce that $\mu=0$ and this is known to the econometrician. In Remark 1, we elaborate further on the effect of estimating the mean.
@@
$l(\theta;y)\propto-\log|\Sigma_{n}(\theta)|-y^{\top}\Sigma_{n}^{-1}(\theta)y,$
and $\Sigma_{n}(\theta)=E_{\theta}\big[(Y_{n}^{\Delta}-\mu)(Y_{n}^{\Delta}-\mu)^{\top}\big]$ is the $n\times n$ covariance matrix of $Y_{n}^{\Delta}$.
The maximum likelihood estimator, $\hat{\theta}_{\text{MLE}}$, is consistent, asymptotically efficient, and follows a limiting normal distribution under standard regularity conditions (e.g., Newey and McFadden, 1994).
However, even in the Gaussian setting, moderate values of $n$ can render numerical optimization of the log-likelihood function infeasible, because the complexity of computing the determinant and inverting the covariance matrix grows very rapidly. For example, deploying an algorithm such as the Cholesky factorization or LU decomposition results in a computational budget of $O(n^{3})$. Less naive approaches exploit the fact that the covariance matrix of a stationary process observed on an equidistant partition has a Toeplitz structure (e.g., Levinson, 1946; Durbin, 1960), yielding a faster calculation speed of $O(n^{2})$. An extra refinement has been achieved with so-called “superfast” algorithms that operate on Toeplitz matrices using the fast Fourier transform (see Brent, Gustavson, and Yun, 1980). The latter exhibit near-linear growth $O(n\log^{2}(n))$ but suffer from numerical instability in practice (see Stewart, 2003). Moreover, the scaling of these algorithms are still inferior to the linear rate $O(n)$. This poses serious issues in a high-frequency setting, where $n$ is often prohibitively large.
Set against this backdrop, we exploit the composite likelihood framework of Lindsay (1988), building on the earlier concept of pseudo-likelihood from Besag (1974) in the spatial setting, see also the survey by Varin, Reid, and Firth (2011). To describe the idea, assume for the moment that $y$ is a sample of length $n$ of a continuous random variable $Y$ defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. [^6] In our setting, the observations are a discrete realization of a continuous-time process. In general, the composite likelihood estimator maximizes a weighted product of likelihoods of marginal or conditional events. We let $f$ denote the density of $Y$ and suppose that $(\mathcal{A}_{1},\dots,\mathcal{A}_{M})$ is a collection of events, $\mathcal{A}_{m}\in\mathcal{F}$, with likelihood $L_{m}(\theta;y)\propto f(y\in\mathcal{A}_{m};\theta)$. The composite likelihood is then defined as
@@
The composite log-likelihood function is, in general, not proportional to the full likelihood, so the model is misspecified (see, e.g., White, 1982). However, the data-generating probability measure is closely related to the assumed parametric distribution, since the marginal (or conditional) densities included in (2.3) are extracted from the true model and correctly specified. Hence, the score of $cl(\theta;y)$ satisfies the first Bartlett identity and forms a collection of unbiased estimating equations. As we show below, it follows by a law of large numbers that $\hat{\theta}_{\mathrm{MCLE}}$ is consistent for the true parameter value, $\theta_{0}$, and, if the memory in the process is not too strong, it is also asymptotically normal. [^8] If the parameterization of the ACF is wrong, $\hat{\theta}_{\mathrm{MCLE}}$ generally does not converge to $\theta_{0}$, but rather a pseudo-parameter, $\theta^{*}$, that minimizes a “composite” Kullback-Leibler divergence between the true and assumed model, i.e. $\theta^{*}=\underset{\theta\in\Theta}{\arg\max}\ \mathbb{E}\left[cl(\theta;Y)\right]$, where the expectation is with respect to the true probability measure.
To showcase the promise of MCLE compared to “brute force” MLE, we simulate an fOU process (defined in Section 3.1) with a true parameter vector given by Panel B in Table 1. The Hurst exponent (the parameter of main interest) is $H=0.1$, so we are in a rough setting. We implement CL as explained in Section 4. In Panel A of Figure 1, we plot the relative runtime of CL versus ML as a function of $n=N\times T$ (measured against the left $y$-axis, while the actual runtime can be read of from the right $y$-axis). Here, $T$ is the number of days in the sample, whereas $N$ is the number of intraday observations. We set $N=12$ and take $T=10,20,\dots,250$. As the sample size grows, CL is completed in a small fraction of the ML runtime. The maximal length of the data vector is n=3,000n=3{,}000 here, which is rather small, but around this point MLE becomes unresponsive (at least on our desktop) or runs out of memory. In Panel B of the figure, we plot the relative root mean squared error (RMSE) of CL versus ML, formatted as before, in terms of estimating $H$. The chasm between the RMSE of MCLE and MLE is relatively constant, because both converge at rate $n^{-1/2}$ in this setting. Hence, here the main drawback of MCLE is the loss of (asymptotic) efficiency.
Figure 1: Comparison of MLE and MCLE.
Panel A: Runtime. Panel B: RMSE.
Refer to caption Refer to caption
Note. We simulate an fOU process, defined in Section 3.1. The three-dimensional parameter vector $\theta=(\kappa,\nu,H)^{\top}$ follows Panel B in Table 1 (the mean is not estimated). The sample size is $n=T\times N$, which means $N$ intraday observations over a period of $T=10,20,\dots,250$ days. As in our simulation and empirical implementation, we fix $N=12$. We estimate $\theta$ with ML and MCL. In Panel A, we show the (actual and relative) runtime, while Panel B shows the associated (actual and relative) root mean squared error (RMSE) of the Hurst parameter, $H$. The true value of $H$ is 0.1.
@@
In the next subsection, we establish an asymptotic theory for a $q$-wise composite likelihood estimator, where the composite likelihood function is formed as an equal weighted product of marginal events defined via the selection of tuples of up to length $q$. It nests both the independence likelihood and the pairwise likelihood as special cases. We take $K\in\mathbb{N}$ fixed and let $Q$ be a collection of $K$ $q^{j}$-tuples of natural numbers, where $q^{j}\in\mathbb{N}$, i.e. tuples of the form $k^{j}=(0,k_{2}^{j},\dots,k_{q^{j}}^{j})\in\{0\}\times\mathbb{N}^{q^{j}-1}$ for $j=1,\dots,K$, indexing which observations to add. Here, we use the convention that $\mathbb{N}^{0}=\emptyset$. To ease notation, we set $q=\max_{j=1,\dots,K}q^{j}$. Moreover, we suppose without loss of generality that the indices are increasing and that $\max_{j=1,\dots,K}k_{q^{j}}^{j}<n$. For $k^{j}\in Q$, we let $f_{k^{j}}(y_{i}^{k^{j}};\theta)=f_{k^{j}}(y_{i},y_{i+k_{2}^{j}},\dots,y_{i+k_{q^{j}}^{j}};\theta)$ denote the density of $Y_{i}^{k^{j}}=(Y_{i\Delta},Y_{(i+k_{2}^{j})\Delta},\dots,Y_{(i+k_{q^{j}}^{j})\Delta})$, which by stationary is independent of ii. [^9] With this notation, maximum likelihood corresponds to $q=n$ and $K=1$.
@@
2.3 Asymptotic theory
To prove our asymptotic theory, we need an identification assumption to ensure that the log-composite likelihood is uniquely maximized at $\theta_{0}$: